OpenStudy (anonymous):

write the equation of the quadratic function with roots 0 and 1 and a vertex at (1/2, -2/3).

OpenStudy (anonymous):

y = a(x - 0) * (x - 1) If x = 1/2 and y = -2/3 -2/3 = a*(1/2)*(1-(1/2) a = -8/3 y = (-8/3) * (x^2-x) y = /8/3x^2 - -8/3x can someone tell me if this is correct

OpenStudy (anonymous):

It should read y = -8/3x^2 -8/3x

OpenStudy (anonymous):

This is correct except you need to continue simplifying: y = -8/3x^2 -8/3x y(3x^2-3x) = -16 3y(x^2 - 1) = -16 y = -16/(3(x-1)(x+1))